# No Black Box Anymore: Demystifying Clinical Predictive Modeling with Temporal-Feature Cross Attention Mechanism

### v2

@@ 2.1 Data
For cohort identification, patients with a confirmed CKD stage 3 diagnosis and complete observation windows were selected. Our prior analysis demonstrated that a 24-month observation window achieves the best modeling performance. Hence, in this study, we focus on the explainability exploration using this 24-month observation window data, comprising 1,422 patients of whom 86 (6%) progressed to ESRD. For this 24-month observation window data, we further divided the period into eight consecutive intervals of three months each, generating structured longitudinal data: the first three months following CKD stage 3 diagnosis were denoted as $t_{0}$, months $3-6$ as $t_{1}$, months 6-9 as $t_{2}$, and continuing similarly up to months $21-24$ labeled as $t_{7}$. Table 1 summarizes the key characteristics of this cohort, highlighting significant differences between patients who progressed to ESRD and those who did not. For comprehensive details on data preparation, feature engineering, and cohort selection, we refer readers to 11.
@@ 2.2 Features
Table 1: Demographic, clinical, and claims-based characteristics of the patient cohort (n=1,422), comparing patients who progressed to ESRD (n=86, 6%) versus those who did not (n=1,336, 94%). Values are presented as mean ± standard deviation for continuous variables and count (percentage) for categorical variables. P-values were calculated using t-tests for continuous variables and chi-squared tests for categorical variables, with significance set at p $<$ 0.05.
@@
Occurrence of CKD stage 4 S4 47 (54.7%) 298 (22.3%) $<$0.001
Occurrence of CKD stage 5 S5 42 (48.8%) 277 (20.7%) $<$0.001
@@ 2.3 Temporal-Feature Cross Attention Mechanism (TFCAM)
We introduce the Temporal-Feature Cross Attention Mechanism (TFCAM), a framework that captures both the temporal dynamics and inter-feature interactions in sequential clinical data. TFCAM leverages three attention layers—temporal, feature-level, and cross-feature—to achieve robust predictions while enhancing interpretability.
@@ 3.2 Temporal-Level Comparison
Figure 1: Temporal attention weights for RETAIN (left) and TFCAM (right). Top row: mean weights across all patients showing U-shaped patterns. Bottom row: attention patterns for ESRD progressors (red) versus non-progressors (blue), with characteristic crossing at t4.
@@ 3.3 Feature-Level Comparison
Both models consistently identified several common high-importance features. Diabetes status received the highest importance score in both models, confirming its established role as a primary discriminating risk factor for CKD progression.’ We saw the same insight in Philipp’s work as well 33. Laboratory values S4 and S5, net expenditure of professionals and outpatient (net_exp_P/O), and hemoglobin levels were ranked within the top five features by both models. Age also appeared among the top six features in both models, reflecting its well-established clinical significance in CKD prognosis.

### v3

@@ 2.1 Data
For cohort identification, patients with a confirmed CKD stage 3 diagnosis and complete observation windows were selected. Our prior analysis demonstrated that a 24-month observation window achieves the best modeling performance. Hence, in this study, we focus on the explainability exploration using this 24-month observation window data, comprising 1,422 patients of whom 86 (6%) progressed to ESRD. For this 24-month observation window data, we further divided the period into eight consecutive intervals of three months each, generating structured longitudinal data: the first three months following CKD stage 3 diagnosis were denoted as $t_{0}$, months 363-6 as $t_{1}$, months 6-9 as $t_{2}$, and continuing similarly up to months 212421-24 labeled as $t_{7}$. Table 1 summarizes the key characteristics of this cohort, highlighting significant differences between patients who progressed to ESRD and those who did not. For comprehensive details on data preparation, feature engineering, and cohort selection, we refer readers to 11.
@@ 2.2 Features
Table 1: Demographic, clinical, and claims-based characteristics of the patient cohort (n=1,422), comparing patients who progressed to ESRD (n=86, 6%) versus those who did not (n=1,336, 94%). Values are presented as mean ± standard deviation for continuous variables and count (percentage) for categorical variables. P-values were calculated using t-tests for continuous variables and chi-squared tests for categorical variables, with significance set at p << 0.05.
@@
Occurrence of CKD stage 4 S4 47 (54.7%) 298 (22.3%) <<0.001
Occurrence of CKD stage 5 S5 42 (48.8%) 277 (20.7%) <<0.001
@@ 2.3 Temporal-Feature Cross Attention Mechanism (TFCAM)
We introduce the Temporal-Feature Cross Attention Mechanism (TFCAM), a framework that captures both the temporal dynamics and inter-feature interactions in sequential clinical data. TFCAM leverages three attention layers — temporal level, feature level, and cross-feature level — to achieve robust predictions while enhancing interpretability.
@@ 3.2 Temporal-Level Comparison
Figure 1: Temporal attention weights for RETAIN (left) and TFCAM (right). Top row: Mean weights across all patients showing U-shaped patterns. Bottom row: Attention patterns for ESRD progressors (red) versus non-progressors (blue), with characteristic crossing at $t_{4}$.
@@ 3.3 Feature-Level Comparison
Both models consistently identified several common high-importance features. Diabetes status received the highest importance score in both models, confirming its established role as a primary discriminating risk factor for CKD progression. We saw the same insight in Philipp’s work as well 33. Laboratory values S4 and S5, net expenditure of professionals and outpatient (net_exp_P/O), and hemoglobin levels were ranked within the top five features by both models. Age also appeared among the top six features in both models, reflecting its well-established clinical significance in CKD prognosis.